OpenStudy (anonymous):

Two corridors whose widths are 64 units and 27 units respectively meet at a right angle. Calculate the length of the longest rigid beam that can pass from one corridor to the other when it is slid along the floor.

OpenStudy (anonymous):

I'm not sure where to even start... we don't have any angles or anything, so...

OpenStudy (amonoconnor):

That's all I can do for you:)

OpenStudy (anonymous):



OpenStudy (anonymous):

The diagram we get looks like that

OpenStudy (anonymous):

@sourwing

OpenStudy (noelgreco):



OpenStudy (noelgreco):

\[\cos a=\frac{ 64 }{ x }\] \[\sin a=\frac{ 27 }{ y }\] Solver for x and y in terms of TRIG a, add, differentiate, and enjoy.